OpenStudy (anonymous):

How many years will it take for a $50 investment to grow to $1000 if it is compounded continuously at a rate of 5%? Round to two decimal places. Do not include units in your answerA(t) = P ert

OpenStudy (anonymous):

this is easy you want 1000 , this is your A ( end amount) you have 50 , this is your P (principal/beginning amount) your rate is 5% , in decimal form , it is .05 , this is R (rate) you solve for T (time) and e is just a special number, about 2.7 A = P e^( R T) 1000 = 50e^(.05 T) divide by 50 20 = e^(.05 T) here you need to use natural logs give a ln to 20 and to e ln (20) = ln e^(.05 T) ln and e cancel out to make 1 ln (20) = .05 T divide by .05 ln 20/.05 = T, or about 59.9

OpenStudy (anonymous):

thank you so much me and my teacher are struggling to find the answer lol thanks !

OpenStudy (anonymous):

59.9146 years

OpenStudy (anonymous):

And sure thing :3